Question Help
How long does it take for an investment to double in value if it is invested at
20​% compounded​ continuously?
At 20% compound continuously, the investment doubles in about __ years.

Answer by rothauserc(4718)

We use the exponential growth or decay formula for continuously compounded interest
:
A = Pe^rt, where P is the original amount, r is the annual interest rate, t is time in years
:
let x be the original amount, then
:
2x = x * e^(0.20t)
:
e^(0.20t) = 2
:
0.20t = ln(2)
:
t = ln(2)/0.20 = 3.4657
:
*************************************************************************
At 20% compound continuously, the investment doubles in about 3.5 years
*************************************************************************
: